# Transition Metal Dichalcogenides Multijunction Solar Cells Toward the Multicolor Limit

### v1

@@ Abstract
Transition metal dichalcogenides (TMDs) and other van der Waals semiconductors enable transfer-printed, lattice-mismatch–free stacking of many photovoltaic junctions, motivating a re-examination of multijunction detailed-balance limits under realistic material and optical constraints. Here we develop an unlimited-junction detailed-balance framework for split-spectrum, multi-terminal vdW multijunction solar cells and apply it to a conservative TMD bandgap window (1.0–2.1 eV). Dynamic-programming optimization shows that the accessible bandgap window imposes a large-$N$ efficiency limit: under full concentration, unconstrained ladders approach 84.5% at $N=50$, whereas the TMD window plateaus near 63.4%. This plateau is set by photons outside the gaps, so radiative quality and optics dominate beyond five junctions for realistic transfer-printed stacks. We identify an experimentally achievable $N=5$ ladder $E_{g}\approx(2.10,1.78,1.50,1.24,1.00)$ eV and map each rung to candidate vdW/TMD absorbers. Using reciprocity and luminescence thermodynamics, we quantify penalties from finite external radiative efficiency, two-sided emission, and luminescent coupling, and we introduce the upward-emitted luminescence power as a computable entropy-loss proxy. Incorporating excitonic absorptance and nanophotonic thickness bounds yields practical thickness and light-management targets for transfer-printed stacks. Finally, inserting an idealized nonreciprocal multijunction model into the reciprocity-optimized ladders provides conservative headroom estimates, consistent with negligible benefit for single junctions but measurable gains for multijunction TMD stacks.
@@
multijunction photovoltaics; transition metal dichalcogenides; detailed balance; external radiative efficiency; luminescent coupling
@@ 1 Introduction
The Shockley–Queisser (SQ) detailed-balance limit for a single-junction solar cell formalizes the thermodynamic tradeoffs between (i) transmission loss from sub-bandgap photons, (ii) thermalization loss from above-bandgap photons, and (iii) radiative emission that reduces the achievable open-circuit voltage $V_{\mathrm{oc}}$.18 Multijunction photovoltaics mitigates these losses by partitioning the solar spectrum across multiple bandgaps.2, 3 In the radiative and ideal-optics limit, the efficiency of multijunction devices approaches the reciprocal multicolor limit as the number of junctions grows.
@@
Despite decades of progress, pushing beyond triple-junction stacks remains challenging in conventional epitaxial material systems because lattice matching, thermal expansion mismatch, and defect formation create practical constraints. Van der Waals semiconductors—especially TMDs—offer a fundamentally different integration pathway: their layered nature enables mechanical exfoliation, deterministic transfer printing, and assembly of many junctions without lattice matching.4 This raises a natural theoretical question: If we could stack an arbitrary number of vdW junctions, and if we could select bandgaps broadly across the solar spectrum, how close could we approach multicolor or even Landsberg-like limits?
@@
Answering this question requires more than bandgap optimization. As emphasized by luminescence thermodynamics, high-efficiency photovoltaics must exhibit strong internal radiative recombination and high external photon extraction.11 Reciprocity further implies that absorption and emission are inseparable: thickness- and nanophotonics-controlled absorptance shapes not only the short-circuit current $J_{\mathrm{sc}}$ but also the radiative dark current and therefore $V_{\mathrm{oc}}$.16, 15 In stacked junctions without intermediate mirrors, an upper junction can emit externally both upward (to the environment) and downward (into the stack), which (i) opens additional optical emission channels and thus introduces an entropy/voltage penalty, and (ii) enables luminescent coupling to lower junctions. These coupling processes can create additional thermalization or entropic losses in the lower junctions, particularly when bandgaps are closely spaced and when the upper junction is highly radiative. Recent detailed-balance analyses of nanophotonic intermediate mirrors in tandem photovoltaics highlight that the net gain can be strongly context dependent, highlighting the need for stack-specific photon-entropy accounting.5
@@
Motivated by this perspective, we pursue a conservative and practical goal: maximize the reciprocal multijunction efficiency achievable with a realistic TMD bandgap window (1.0–2.1 eV), and quantify how luminescence extraction and thickness constraints shape the approach to the multicolor limit. Nonreciprocal optics is treated as an optional strategy: we benchmark its theoretical headroom by applying the idealized nonreciprocal multijunction model of Fan et al.13 to the same bandgap ladders optimized under reciprocity (a conservative “plug-in” comparison), and we explicitly contrast this multijunction headroom with the absence of benefit in single-junction photovoltaics.12
Beyond bandgap engineering, system-level photonics layers can further reshape the spectral and angular photon budget in practical photovoltaics. Examples include luminescent solar concentrators for spectral management,9 neutral-colored transparent crystalline-silicon photovoltaics,8 and transparent radiative-cooling exterior films for outdoor deployment.1 These developments reinforce the motivation for a unified thermodynamic treatment of emission, coupling, and thickness constraints when designing transfer-printed multijunction stacks.
2 Overview of the device concept and photon-flow constraints
@@
Figure 1 sketches the transfer-printed vdW multijunction architecture considered here: a stack of independently contacted subcells assembled by vdW stacking or transfer printing, with selective contacts/interlayers between subcells, a transparent top electrode, and a reflective back electrode. Throughout, we focus on multi-terminal extraction (each subcell operates at its own maximum power point), which isolates the fundamental spectral-partitioning thermodynamics from current-matching constraints.
@@
Figure 1: Conceptual schematic of a transfer-printed vdW/TMD multijunction solar cell (representative $N=5$ ladder). In a reciprocal stack without intermediate mirrors, external luminescence from an upper junction can exit upward ($\Phi^{\uparrow}_{\mathrm{ext}}$) and propagate downward ($\Phi^{\downarrow}_{\mathrm{ext}}$), enabling luminescent coupling to lower junctions and opening additional emission channels (entropy/voltage penalty). Intermediate mirrors or angular filters can suppress unwanted emission channels and recover higher radiative voltage.11, 15 An optional nonreciprocal optical layer can, in principle, further reshape emission pathways and access additional thermodynamic headroom in multijunction stacks.13
@@
Two points are central. First, in the detailed-balance limit each junction is a light-emitting diode under forward bias: at open circuit and near maximum power, the device emits a substantial photon flux. Strong external luminescence is therefore not a loss to be suppressed, but a signature of a high-quality device approaching the SQ limit.11 Second, in a stacked architecture, the direction and angular distribution of luminescence matter. If an upper junction emits downward into the stack, that emission is not necessarily “wasted”—it can be re-absorbed by lower junctions (luminescent coupling). However, it can still introduce (i) a voltage penalty for the emitter by opening emission channels, and (ii) a thermalization penalty in the absorber if the photon energy significantly exceeds the lower-junction bandgap. A quantitative framework must therefore track both photon flux and power flux of luminescence.
@@ 3.1 Spectral model, concentration, and incident power
$\Phi_{\mathrm{bb}}(E,T)=\frac{2E^{2}}{h^{3}c^{2}}\,\frac{1}{\exp(E/k_{B}T)-1},$ (1)
where $E$ is photon energy, $h$ is Planck’s constant, and $c$ is the speed of light. The sun subtends solid angle $\Omega_{s}=\pi\sin^{2}\theta_{s}$ where $\theta_{s}\simeq$0.266\text{\,}\mathrm{d}\mathrm{e}\mathrm{g}\mathrm{r}\mathrm{e}\mathrm{e}$\$ is the solar half-angle. Optical concentration is represented by scaling the incident flux by a factor $C$, so that the spectral photon flux incident on the device is
$\Phi_{\odot}(E)=C\,\Omega_{s}\,\Phi_{\mathrm{bb}}(E,T_{s}).$ (2)
@@
$P_{\mathrm{in}}(C)=\int_{0}^{\infty}E\,\Phi_{\odot}(E)\,dE.$ (3)
@@ 3.2 Single-junction detailed balance
$J(V)=J_{\mathrm{sc}}-J_{0}\left[\exp\left(\frac{qV}{k_{B}T_{c}}\right)-1\right],$ (4)
where $J_{\mathrm{sc}}$ is the short-circuit current density and $J_{0}$ is the dark-current prefactor. For an absorber with absorptance $A(E)$ (not necessarily unity), the photogenerated current density is
$J_{\mathrm{sc}}=q\int_{0}^{\infty}A(E)\,\Phi_{\odot}(E)\,dE.$ (5)
Radiative recombination produces an emitted photon flux proportional to $A(E)\,\Phi_{\mathrm{bb}}(E,T_{c})$ by Kirchhoff’s law and reciprocity.16 Thus the radiative dark current is
$J_{0,\mathrm{rad}}=q\int_{0}^{\infty}A(E)\,\Phi_{\mathrm{bb}}(E,T_{c})\,\Omega_{\mathrm{emit}}\,dE,$ (6)
@@
Nonradiative recombination and parasitic optical loss reduce the fraction of recombination events yielding externally emitted photons. A convenient device-quality metric is the external radiative efficiency $\mathrm{ERE}$, which can be introduced via
$J_{0}=\frac{J_{0,\mathrm{rad}}}{\mathrm{ERE}}.$ (7)
@@
$\Delta V_{\mathrm{ERE}}=\frac{k_{B}T_{c}}{q}\ln(\mathrm{ERE}),$ (8)
identical in form to the $\ln Q^{\mathrm{LED}}_{e}$ term that appears in the thermodynamic voltage decomposition of Rau and co-workers.15, 14, 7
@@
$v_{\mathrm{mpp}}=W\!\left(e\left[1+J_{\mathrm{sc}}/J_{0}\right]\right)-1,$ (9)
with $V_{\mathrm{mpp}}=(k_{B}T_{c}/q)\,v_{\mathrm{mpp}}$ and $P_{\mathrm{mpp}}=V_{\mathrm{mpp}}J(V_{\mathrm{mpp}})$. A full derivation is provided in the Supporting Information.
@@ 3.3 Unlimited-junction split-spectrum multijunction model
We consider an $N$-junction multijunction device with bandgaps $E_{g,1}>E_{g,2}>\cdots>E_{g,N}$. In the split-spectrum/multi-terminal limit, junction $i$ interacts with photon energies in a window $E\in[E_{g,i},E_{g,i-1})$ where $E_{g,0}\equiv E_{\mathrm{max}}$ is a numerical high-energy cutoff. For step absorbers, $A_{i}(E)=1$ in this window and $A_{i}(E)=0$ outside it. The total maximum power density is
$P^{\ast}_{\mathrm{tot}}=\sum_{i=1}^{N}P_{i}^{\ast}(E_{g,i};E_{g,i-1}),$ (10)
where $P_{i}^{\ast}$ is obtained from Eqs. (5)–(9) restricted to the window. We compute optimal bandgap ladders by maximizing Eq. (10) with dynamic programming (DP) on a discretized bandgap grid (see Supporting Information). We report two scenarios: (i) unconstrained $E_{g}$ (theoretical convergence), and (ii) a conservative, experimentally motivated TMD window $E_{g}\in[1.0,2.1]$ eV.
@@ 3.4 Reciprocity, emission channels, and luminescent coupling in reciprocal stacks
The reciprocity relation between photovoltaic quantum efficiency and electroluminescent emission connects absorption and emission spectra.16 A compact form for the emitted photon flux spectrum under bias $V$ is
$\phi_{\mathrm{em}}(E,V)=A(E)\,\Phi_{\mathrm{bb}}(E,T_{c})\,\Omega_{\mathrm{emit}}\left[\exp\left(\frac{qV}{k_{B}T_{c}}\right)-1\right],$ (11)
where $A(E)$ is the device absorptance (or external PV quantum efficiency under ideal collection). Equation (11) makes explicit that any thickness- or nanophotonics-induced change in $A(E)$ affects both $J_{\mathrm{sc}}$ and $J_{0,\mathrm{rad}}$.
In a reciprocal multijunction stack lacking intermediate mirrors, each junction can emit both upward and downward (two hemispheres). Allowing both hemispheres doubles the radiative dark current prefactor and produces a radiative-voltage reduction $\Delta V=(k_{B}T_{c}/q)\ln 2$ at fixed generation. More importantly, downward emission from junction $i$ can be absorbed by junction $i+1$ if $E_{g,i}>E_{g,i+1}$, producing a luminescent-coupling current $J_{\mathrm{LC}}$. We model this effect using a minimal coupled detailed-balance chain (see Supporting Information) in which
$J_{i}(V_{i})$ $=J_{\odot,i}+J^{\downarrow}_{\mathrm{rad},i-1}(V_{i-1})-\frac{J_{0,i}^{\uparrow}+J_{0,i}^{\downarrow}}{\mathrm{ERE}_{i}}\left[\exp\left(\frac{qV_{i}}{k_{B}T_{c}}\right)-1\right],$ (12)
$J^{\downarrow}_{\mathrm{rad},i}(V_{i})$ $=J_{0,i}^{\downarrow}\left[\exp\left(\frac{qV_{i}}{k_{B}T_{c}}\right)-1\right],$ (13)
with $J^{\downarrow}_{\mathrm{rad},0}\equiv 0$. This model captures (i) the emission-channel entropy penalty (two-sided emission increases $J_{0}$), and (ii) the photon recycling/coupling that can boost the lower-junction current. We solve for the system maximum power point by coordinate-ascent optimization over the set of subcell voltages $\{V_{i}\}$.

### v3

@@ Abstract
Transition metal dichalcogenides (TMDs) and other van der Waals (vdW) semiconductors enable transfer-printed, lattice–mismatch-free stacking of many photovoltaic junctions, motivating a re-examination of multijunction detailed-balance limits under realistic material and optical constraints. Here, we develop a unified thermodynamic framework for a multijunction photovoltaic device, which can define a clear set of device-window constraints, optical boundary conditions, and luminescence/entropy penalties and therefore define how closely any realistic multijunction photovoltaic device can approach multicolor limit. By applying it to a conservative TMD bandgap window (1.0–2.1 eV), we show that the accessible bandgap window imposes a large-junction number ($N$) efficiency limit: under full concentration, unconstrained ladders approach 84.5% at $N=50$, whereas the TMD window plateaus near 63.4%. This efficiency plateau is set by photons outside the bandgaps, so radiative quality and optics dominate beyond $N=5$ junctions for realistic transfer-printed device stacks. We identify an experimentally achievable $N=5$ ladder $E_{g}\approx(2.10,1.78,1.50,1.24,1.00)$ eV and map each rung to candidate vdW/TMD absorbers. Using reciprocity and luminescence thermodynamics, we quantify penalties from finite external radiative efficiency, two-sided emission, and luminescent coupling, and introduce the upward-emitted luminescence power as an indicator of entropy-loss proxy. Incorporating excitonic absorptance and nanophotonic thickness bounds yields practical thickness and light-management targets for transfer-printed stacks. Finally, inserting an idealized nonreciprocal multijunction model into the reciprocity-optimized ladders provides conservative efficiency advantage estimates, which are consistent with negligible benefit for single junctions but measurable efficiency gains for multijunction TMD devices.
@@
multijunction photovoltaics , transition metal dichalcogenides , detailed balance , external radiative efficiency , luminescent coupling
@@ 1 Introduction
The Shockley–Queisser (SQ) detailed-balance limit for a single-junction solar cell formalizes the thermodynamic tradeoffs between (i) transmission loss from sub-bandgap photons, (ii) thermalization loss from above-bandgap photons, and (iii) radiative emission that reduces the achievable open-circuit voltage $V_{\mathrm{oc}}$.[16] Multijunction photovoltaics mitigates these losses by partitioning the solar spectrum across multiple bandgaps.[1, 2, 4] In the radiative and ideal-optics limit, the efficiency of multijunction devices approaches the reciprocal multicolor limit as the number of junctions ($N$) grows.
@@
Despite decades of progress, pushing beyond triple-junction stacks remains a challenge in conventional epitaxial material systems, because lattice matching, thermal expansion mismatch, and defect formation lead to practical constraints. Van der Waals (vdWs) semiconductors, especially transition metal dichalcogenides (TMDs), offer a fundamentally different integration pathway: their layered nature enables mechanical exfoliation, deterministic transfer printing, and assembly of many junctions without lattice matching.[3, 18, 15] This raises a natural theoretical question: if we could stack an arbitrary number of vdW junctions and could select bandgaps broadly across the solar spectrum, how close could we approach multicolor or even Landsberg limits?
@@
Answering this question requires more than bandgap optimization. As emphasized by luminescence thermodynamics, high-efficiency photovoltaics must exhibit strong internal radiative recombination and high external photon extraction.[9, 5, 6, 14, 7] Reciprocity further implies that absorption and emission are inseparable: thickness- and nanophotonics-controlled absorptance defines not only the short-circuit current $J_{\mathrm{sc}}$ but also the radiative dark current and therefore $V_{\mathrm{oc}}$.[14, 13] In stacked junctions without intermediate mirrors,[4] an upper junction can emit externally both upward (to the environment) and downward (into the stack), which (i) opens additional optical emission channels and thus introduces an entropy and voltage penalty, and (ii) enables luminescent coupling to lower junctions.[1, 2, 4] These coupling processes can give rise to additional thermalization or entropic losses in the lower junctions, particularly when the bandgaps are closely spaced and the upper junction is highly radiative. Recent detailed-balance analyses of nanophotonic intermediate mirrors in tandem photovoltaics highlight that the net gain can be strongly device configuration-dependent, highlighting the need for stack-specific photon-entropy accounting.[4]
@@
Motivated by this perspective, we pursue a conservative and practical goal: maximize the reciprocal multijunction efficiency achievable with a realistic TMD bandgap window (1.0–2.1 eV), and quantify how both luminescence extraction and thickness constraints define the approach to the multicolor limit. Nonreciprocal optics is treated as an optional strategy: we benchmark its theoretical efficiency gain by applying the idealized nonreciprocal multijunction model of Fan et al.[11] to the same bandgap ladders, which are optimized under reciprocity (a conservative “plug-in” comparison). Then, we systematically contrast this multijunction efficiency gain with the absence of benefit in single-junction photovoltaics.[10] These developments reinforce the motivation for a unified thermodynamic treatment of emission, coupling, and thickness constraints when designing transfer-printed TMD multijunction stacks.
2 The device concept and photon-flow constraints
@@
Figure 1 summarizes the transfer-printed TMD multijunction architecture, studied in this work: a stack of independently contacted subcells assembled by TMD stacking or transfer printing, with selective contacts/interlayers between subcells, a transparent top electrode, and a reflective back electrode. Throughout, we focus on multi-terminal extraction (each TMD subcell operates at its own maximum power point), which isolates the fundamental spectral-partitioning thermodynamics from current-matching constraints.
@@
Figure 1: Conceptual schematic of a transfer-printed vdW/TMD multijunction solar cell (representative $N=5$ ladder). In a reciprocal stack without intermediate mirrors, external luminescence from an upper junction can exit upward ($\Phi^{\uparrow}_{\mathrm{ext}}$) and propagate downward ($\Phi^{\downarrow}_{\mathrm{ext}}$), enabling luminescent coupling to lower junctions and opening additional emission channels (entropy/voltage penalty). Intermediate mirrors or angular filters can suppress unwanted emission channels and recover higher radiative voltage.[9, 13] An optional nonreciprocal optical layer can, in principle, further reshape emission pathways and access additional thermodynamic efficiency gain in multijunction stacks.[11]
@@
Two points are central. First, in the detailed-balance limit, each junction is a light-emitting diode under forward bias: at open circuit and near maximum power, the device emits a substantial photon flux. Strong external luminescence is therefore not a loss to be suppressed, but a signature of a high-quality device approaching the SQ limit.[9] Second, in a stacked architecture, the direction and angular distribution of luminescence matter. If an upper junction emits downward into the stack, that emission is not necessarily “wasted.” It can be re-absorbed by lower junctions (luminescent coupling). However, it can still introduce (i) a voltage penalty for the emitter by opening emission channels, and (ii) a thermalization penalty in the absorber if the photon energy significantly exceeds the lower-junction bandgap. A quantitative framework must therefore track both photon flux and power flux of luminescence.
@@ 3.1 Spectral model, concentration, and incident power
$\Phi_{\mathrm{bb}}(E,T)=\frac{2E^{2}}{h^{3}c^{2}}\,\frac{1}{\exp(E/k_{B}T)-1}$ (1)
where $E$ is photon energy, $h$ is Planck’s constant, and $c$ is the speed of light. The Sun subtends solid angle $\Omega_{s}=\pi\sin^{2}\theta_{s}$, where $\theta_{s}\simeq$0.266\text{\,}\mathrm{d}\mathrm{e}\mathrm{g}\mathrm{r}\mathrm{e}\mathrm{e}$\$ is the solar half-angle. Optical concentration is represented by scaling the incident flux by a factor $C$, so that the spectral photon flux, which is incident on the device, is
$\Phi_{\odot}(E)=C\,\Omega_{s}\,\Phi_{\mathrm{bb}}(E,T_{s})$ (2)
@@
$P_{\mathrm{in}}(C)=\int_{0}^{\infty}E\,\Phi_{\odot}(E)\,dE$ (3)
@@ 3.2 Single-junction detailed balance
$J(V)=J_{\mathrm{sc}}-J_{0}\left[\exp\left(\frac{qV}{k_{B}T_{c}}\right)-1\right]$ (4)
where $J_{\mathrm{sc}}$ and $J_{0}$ are the short-circuit current density and the dark-current prefactor, respectively. For an absorber with absorptance $A(E)$ (not necessarily unity), the photogenerated current density is
$J_{\mathrm{sc}}=q\int_{0}^{\infty}A(E)\,\Phi_{\odot}(E)\,dE$ (5)
Radiative recombination produces an emitted photon flux proportional to $A(E)\,\Phi_{\mathrm{bb}}(E,T_{c})$ by both Kirchhoff’s law and reciprocity.[14] Thus, the radiative dark current is
$J_{0,\mathrm{rad}}=q\int_{0}^{\infty}A(E)\,\Phi_{\mathrm{bb}}(E,T_{c})\,\Omega_{\mathrm{emit}}\,dE$ (6)
@@
Nonradiative recombination and parasitic optical loss reduce the fraction of recombination events, which are responsible for externally emitted photons. A convenient device-quality metric is the external radiative efficiency ($\mathrm{ERE}$), which can be introduced via
$J_{0}=\frac{J_{0,\mathrm{rad}}}{\mathrm{ERE}}$ (7)
@@
$\Delta V_{\mathrm{ERE}}=\frac{k_{B}T_{c}}{q}\ln(\mathrm{ERE})$ (8)
which is identical in form to the $\ln Q^{\mathrm{LED}}_{e}$ term that appears in the thermodynamic voltage decomposition of Rau and co-workers.[13, 12, 6]
@@
$v_{\mathrm{mpp}}=W\!\left(e\left[1+J_{\mathrm{sc}}/J_{0}\right]\right)-1$ (9)
with $V_{\mathrm{mpp}}=(k_{B}T_{c}/q)\,v_{\mathrm{mpp}}$ and $P_{\mathrm{mpp}}=V_{\mathrm{mpp}}J(V_{\mathrm{mpp}})$. A full derivation is provided in the Supplement Information.
@@ 3.3 Unlimited-junction split-spectrum multijunction model
We consider a $N$-junction multijunction device with bandgaps $E_{g,1}>E_{g,2}>\cdots>E_{g,N}$. In the split-spectrum and multi-terminal limit, junction $i$ interacts with photon energies in a window $E\in[E_{g,i},E_{g,i-1})$, where $E_{g,0}\equiv E_{\mathrm{max}}$ is a numerical high-energy cutoff. For step absorbers, $A_{i}(E)=1$ in this window and $A_{i}(E)=0$ outside it. The total maximum power density is
$P^{\ast}_{\mathrm{tot}}=\sum_{i=1}^{N}P_{i}^{\ast}(E_{g,i};E_{g,i-1})$ (10)
where $P_{i}^{\ast}$ is obtained from Eqs. (5)–(9), which are restricted to the window. We compute optimal bandgap ladders by maximizing Eq. (10) with dynamic programming (DP) on a discretized bandgap grid (see Supplement Information). We study two scenarios: (i) unconstrained $E_{g}$ (theoretical convergence) and (ii) a conservative, experimentally available TMD bandgap window $E_{g}\in[1.0,2.1]$ eV.
@@ 3.4 Reciprocity, emission channels, and luminescent coupling in reciprocal stacks
The reciprocity relation between photovoltaic quantum efficiency and electroluminescent emission connects absorption and emission spectra.[14] A compact form for the emitted photon flux spectrum under bias $V$ is
$\phi_{\mathrm{em}}(E,V)=A(E)\,\Phi_{\mathrm{bb}}(E,T_{c})\,\Omega_{\mathrm{emit}}\left[\exp\left(\frac{qV}{k_{B}T_{c}}\right)-1\right]$ (11)
where $A(E)$ is the device absorptance (or external quantum efficiency under ideal collection). Equation (11) implies that any thickness- or nanophotonics-induced change in $A(E)$ affects both $J_{\mathrm{sc}}$ and $J_{0,\mathrm{rad}}$.
In a reciprocal multijunction stack lacking intermediate mirrors, each junction can emit both upward and downward (two hemispheres). Allowing both forward and downward hemispheres doubles the radiative dark current prefactor and produces a radiative-voltage reduction $\Delta V=(k_{B}T_{c}/q)\ln 2$ at fixed generation. More importantly, downward emission from junction $i$ can be absorbed by junction $i+1$, if $E_{g,i}>E_{g,i+1}$, producing a luminescent-coupling current $J_{\mathrm{LC}}$. We model this effect using a minimal coupled detailed-balance chain (see Supplement Information), in which
$J_{i}(V_{i})$ $=J_{\odot,i}+J^{\downarrow}_{\mathrm{rad},i-1}(V_{i-1})-\frac{J_{0,i}^{\uparrow}+J_{0,i}^{\downarrow}}{\mathrm{ERE}_{i}}\left[\exp\left(\frac{qV_{i}}{k_{B}T_{c}}\right)-1\right]$ (12)
$J^{\downarrow}_{\mathrm{rad},i}(V_{i})$ $=J_{0,i}^{\downarrow}\left[\exp\left(\frac{qV_{i}}{k_{B}T_{c}}\right)-1\right]$ (13)
with $J^{\downarrow}_{\mathrm{rad},0}\equiv 0$. This model captures (i) the emission-channel entropy penalty (i.e., two-sided emission increases $J_{0}$), and (ii) the photon recycling and coupling that can boost the lower-junction current. We solve for the system maximum power point by coordinate-ascent optimization over the set of subcell voltages $\{V_{i}\}$.